OpenStudy (anonymous):

1.5q-1.4-1.3q=-0.8q-3.9

hero (hero):

q = -2.6

OpenStudy (anonymous):

i got 2.5 could you show your work love ?

OpenStudy (akshay_budhkar):

(1.5-1.3)q+0.8q=1.4-3.9 therefore q=-2.5

OpenStudy (anonymous):

2.5 is correct .

hero (hero):

I did it in my head so there was no work to show, but I goofed up by 0.1

OpenStudy (akshay_budhkar):

Hero you r solving maths all day a silly mistake is acceptable man no worries lol

OpenStudy (anonymous):

- Agreed.. ;))